# Prompt-Guided Latent Diffusion with Predictive Class Conditioning for 3D Prostate MRI Generation

### v1

@@ Abstract
Latent diffusion models (LDM) could alleviate data scarcity challenges affecting machine learning development for medical imaging. However, medical LDM training typically relies on performance- or scientific accessibility-limiting strategies including a reliance on short-prompt text encoders, the reuse of non-medical LDMs, or a requirement for fine-tuning with large data volumes. We propose a Class-Conditioned Efficient Large Language model Adapter (CCELLA) to address these limitations. CCELLA is a novel dual-head conditioning approach that simultaneously conditions the LDM U-Net with non-medical large language model-encoded text features through cross-attention and with pathology classification through the timestep embedding. We also propose a joint loss function and a data-efficient LDM training framework. In combination, these strategies enable pathology-conditioned LDM training for high-quality medical image synthesis given limited data volume and human data annotation, improving LDM performance and scientific accessibility. Our method achieves a 3D FID score of 0.025 on a size-limited prostate MRI dataset, significantly outperforming a recent foundation model with FID 0.071. When training a classifier for prostate cancer prediction, adding synthetic images generated by our method to the training dataset improves classifier accuracy from 69% to 74%. Training a classifier solely on our method’s synthetic images achieved comparable performance to training on real images alone.
@@ 1 Introduction
Machine learning (ML) for medical image interpretation is a rapidly expanding field intent on improving imaging-based healthcare. While ML continues to improve outside of medical imaging, recent ML advancements require training data in higher orders of magnitude than their predecessors. These advances cannot translate easily to the medical imaging space due to labeled data scarcity challenges including patient privacy and the cost of expert annotation. Generative approaches, including latent diffusion models (LDM), provide a possible solution to labeled training data scarcity, whereby synthetic data could be used to artificially increase ML training data volume [1].
@@
Developing LDMs for medical imaging poses a few challenges. Training LDMs from scratch require large volumes of data. The Stable Diffusion LDM [2] for generating nonmedical, 2D color images was trained initially on 400 million image-text pairs [3], with subsequent versions training on 5.85 billion image-text pairs [4]. These data volumes are inaccessible to most healthcare institutions due to aforementioned data scarcity challenges. One approach is to fine-tune non-medical (2D) LDMs on 2D medical imaging data [5], however this strategy cannot easily extend to 3D. As an added challenge, because 3D medical images typically have more voxels than 2D images have pixels, 3D medical image generation should theoretically require more training data. Another approach is to add labeled data conditioning to increase data efficiency during training [6]. While some medical LDMs have attempted to use the pre-trained CLIP text encoder to condition the LDM, similar to Stable Diffusion [2], research has shown that the CLIP text encoder mostly emphasizes the first 20 tokens [7]. Thus, rather than using detail-rich clinical reports, medical imaging LDM studies typically use only short, curated medical phrases or keywords with either the existing [8] or fine-tuned [5] CLIP text encoder. These strategies require extra annotation and preprocessing or volumes of data that are challenging for a single institution to gather, limiting the scientific accessibility of medical LDM development.
@@
In this study we propose Class-Conditioned Efficient Large Language Model Adapter (CCELLA), a text adapter extending ELLA [9], to improve LDM performance for synthetic medical image generation in a challenging medical imaging domain. CCELLA processes the text embedding generated by a large language model (LLM) from clinical reports prior to dual-head conditioning of the LDM U-Net. CCELLA performs two functions simultaneously: the first head aligns the text embedding with the image latent space prior to conditioning; the second head provides class conditioning for the LDM during training. Each of these functions have independently demonstrated improvements in LDM performance outside of medical imaging [9, 10]. This joint function is hypothesized to provide two advantages: 1) to assert that the LDM receives conditioning about the presence or absence of pathology within the image, 2) more importantly, to inform CCELLA which text features should be emphasized when aligning the embedding to the image latent space for LDM conditioning. To achieve these functions, we also propose a modified loss function that simultaneously trains the LDM to produce high quality medical images and to appropriately extract pathology information from the clinical report text.
@@
Finally, we propose an LDM training framework around CCELLA that uses limited human data annotation and a training data volume reasonably accessible to a single healthcare institution, with the intent of improving the scientific accessibility of LDMs for medical imaging despite limited available data volumes. The framework functions by exploiting the clinical report and pathology labeling already performed by radiologists during clinical routine and by reducing training requirements with publicly available, pre-trained components [1, 11].
@@
This study focused specifically on MRI for prostate cancer as the target domain to evaluate this framework. This pathology and imaging method was selected due to its difficulty in clinical interpretation, its moderate inter-reader variability between clinical experts, and the relatively low-availability of publicly available domain-specific data with the largest dataset consisting of 1500 samples [12, 13]. This study then performed a series of ablations to assess the mechanism by which CCELLA improves LDM performance and compared performance differences among variants of the proposed adapter. Finally, to validate utility of the proposed LDM training framework for dataset augmentation, this study demonstrated that the synthetic data generated with our proposed LDM framework can improve a classifier model at interpreting prostate MRI for cancer presence.
@@
  • A novel dual-head adapter for simultaneous alignment of medical text features extracted by a non-medical LLM to the medical imaging space and pathology class extraction from medical text
  • A LDM training framework for improved pathology-conditioned synthetic medical image generation that is trainable with a reasonable data volume.
@@ 2.1 Latent Diffusion Models
2.2 Latent Diffusion Models within Medical Imaging
LDMs have shown promising results in generating synthetic images both outside and within medical imaging [16]. Because LDM research focuses primarily on 2D, non-medical color image generation on large datasets, medical imaging LDM studies typically also require large data volumes or strategic data-efficient methods. For example, Roentgen [5] adapted non-medical LDMs for chest x-ray generation by fine-tuning the Stable Diffusion LDM [2] on a public dataset with over 350k images. Conversely, de Wilde et al. froze Stable Diffusion and trained specific text prompts to generate 2D prostate MRI slices with or without suspicious lesions [17].
Synthetic medical image generation in 3D is faced with additional challenges related to the increased number of pixels per image compared to 2D LDMs. To combat LDM hallucinations in generating 3D CT images, Xu et al. split the synthetic image generation process into low- and high-resolution stages while generating both images and segmentation maps simultaneously in the low-resolution stage [18]. The upscaled segmentation maps then enabled higher-quality synthetic image generation in the high-resolution LDM stage. MAISI focused instead on computational efficiency and developed a secondary LDM to augment real segmentation masks for use in generating synthetic 3D CT images [1].
2.3 Image Conditioning with Diffusion Models
@@
Conditioning allows for control of the denoising U-Net behavior during LDM training and inference. LDM conditioning typically takes up to three forms. The first is timestep embedding, informing the LDM of its stage in the denoising process. Similar to positional encoding in transformers [19], the timestep is encoded with sinusoidal encoding and then conditions the LDM through addition with the latent image. The second is prompt embedding, informing the LDM of additional information during denoising. Prompt embedding can take different forms but typically are images or text. Prompting data are encoded with separate encoders and condition the LDM through concatenation with the latent image (usually for image prompting) or by cross-attention at multiple layers within the denoising U-Net (usually for text prompting). Class conditioning is a lesser-used third method that expands upon the timestep embedding by adding or concatenating a vector with class information to the timestep embedding [10]. This allows for class-based information to condition the U-Net during denoising. For example, MAISI was trained to accept segmentation masks as prompt embedding, and included both image voxel spacing and body region as class conditioning that modified the timestep embedding [1].
@@
LDM text conditioning typically relies on the Contrastive Language-Image Pre-Training (CLIP) text encoder [20], trained to encode text to a shared image-text latent space. Beyond being trained on non-medical text, the CLIP text encoder is limited by its 77 text token input restriction, with literature documenting an effective length of under 20 tokens [21]. Consequently, text prompting for commercial LDMs are typically restricted to keywords or short phrases. Within medicine however, clinical reports are typically longer and rich in free-text information including location, disease severity, or other defining characteristics that could improve the realism of synthetic image generation [6].
Strategies for text prompting in medical imaging LDMs include using short phrases [22], training new embeddings for the frozen CLIP text encoder [17], fine-tuning the CLIP text encoder on medical data [5], or training new LDMs with a domain-relevant medical LLM [5, 18].
@@
PathLDM is the most similar in motivation to our proposed work [6]. Recognizing the richness of clinical text, the value in using class-based information, the utility of automated ground-truth labeling, and the limitations to the CLIP text encoder for conditioning LDMs, PathLDM developed an automated text prompt generation workflow for synthetic 2D breast cancer histopathology slide generation. This workflow extracted two pathology-related probabilities (“Tumor” and “TIL”) with pretrained models that were each assigned “high” or “low” after probability thresholding, and summarized each slide’s clinical report to under 75 words using the GPT-3.5 LLM [24]. PathLDM also replaced the CLIP text encoder with a pretrained, histopathology-specific CLIP text encoder variant [25]. The final text prompt for LDM conditioning consisted of two class-based keywords and the summarized report in the form “[high/low] Tumor, [high/low] TIL, [report summary]”, allowing PathLDM to receive both classifier and key text report information as conditioning.
@@
  • Uses text from the original report instead of a LLM-generated summary;
  • Trains the class extractor simultaneously with the LDM, rather than relying on a pretrained solution for class extraction;
@@
  • Performs class conditioning using timestep embedding rather than through text encoding.
@@
  • Modifies each TSC block by adding pathology class extractor layers;
@@ 3.1 Dataset
A hold-out test set of 740 images was randomly selected on a per-case level and used to assess the results in this study. Clinical reports were available for 4150 studies from the healthcare institution. No public dataset used in this study included full clinical reports with their images. Because some images from the healthcare institution were acquired prior to the latest update in prostate MRI reporting standard, these cases were reviewed and assigned a new pathology classification per the updated reporting guidelines [27]. All patients had either a radiology classification or histopathology results from within 6 months of the MRI. In the event a radiology classification was not present for an image within the training set, the histopathology results were used instead. Images were center-cropped to a $128mm\times 128mm$ field of view in the axial plane, normalized and then resampled to $256\times 256\times 128$ voxel dimensions using Lanczos interpolation [28].
3.2 Implementation
The proposed framework in this study uses two pretrained components. The first is the MAISI LDM VQ-VAE, pretrained on 3D CT and MRI images [1]. The second is Google’s FLAN-T5 XXL LLM [11]. The base LDM in this study used the frozen MAISI VQ-VAE with a learnable 4-channel, 4-layer denoising U-Net that concatenated image voxel spacings to the timestep embedding. The frozen Flan-T5 XXL text encoder was used to encode the clinical report text.
CCELLA extends the ELLA adapter from [23] by adapting the encoded text for both deep cross-attention and timestep embedding conditioning. CCELLA begins with six class-conditioning TSC (CC-TSC) blocks, each consisting of a TSC block with a separate classifier head of two fully-connected layers of hidden size 256 and output size 2 (MLP block) at the end of the block. The outputs from all CC-TSC classifier heads were concatenated and fed into a final fully-connected layer to generate an output of size 2 to be concatenated with the timestep embedding (Figure 1).
Figure 1: Overview of CCELLA. Consists of six CC-TSC blocks with classifier and aligned text embedding heads. TSC block description can be found in [9].
@@ 3.2 Implementation
The intent of this loss function is to emphasize both LDM image reconstruction as well as appropriate pathology extraction from the reference text. For this study, we used L1 loss for reconstruction and Focal Loss for class prediction, with class loss weighting $\lambda=10^{-4}$.
3.3 Training and Assessment
CCELLA design can be found in Figure 1, providing both cross-attention and timestep conditioning to the LDM. CCELLA performance was assessed against three baseline LDMs: 1) MAISI without text or class conditioning (BASE), 2) MAISI with radiology classification timestep conditioning (RADBASE), 3) MAISI with text conditioning using FLAN-T5 XXL and a chain of six TSC blocks (ELLABASE). All LDMs were trained for 400 epochs with initial learning rate of $10^{-4}$, the AdamW optimizer, a second-order polynomial learning rate scheduler, and batch size of 16 on four Tesla V100 GPUs. All models received image voxel spacing as timestep conditioning as per the foundation MAISI model [1]. During the training of LDMs with text conditioning, images that did not have a text report used an empty string as a proxy text report. When training CCELLA, $\lambda$ in (1) was set to 0 for these cases so that the classifier portion of the loss function did not update the model for these images.
Fréchet Inception Distance (FID) was used to evaluate LDM performances [29]. FID was calculated in two ways: 1) in 3-dimensions using the Med3D ResNet50 pretrained on 23 datasets [30], 2) in 2-dimensions for each of the sagittal, coronal and axial planes using a ResNet50 pretrained on RadImageNet [31].
3.4 Ablation Studies
Two ablation studies were performed. The first assessed the components of CCELLA, namely the joint conditioning and classifier heads and the joint loss function. This ablation study compared CCELLA to 1) an LDM with ELLA text conditioning and separate timestep conditioning using ground truth radiology classification (DECOUPLED), 2) CCELLA using only L1 loss instead of the modified loss in (1) (L1ONLY). The second ablation study compared CCELLA to variants by: 1) placing only one CC-TSC block after a chain of five TSC blocks (LAST), 2) placing one CC-TSC block before five TSC blocks (FIRST), 3) chaining six CC-TSC blocks where class adapter weights are shared across all blocks (SHARED).

### v2

@@ Abstract
Objective: Latent diffusion models (LDM) could alleviate data scarcity challenges affecting machine learning development for medical imaging. However, medical LDM strategies typically rely on short-prompt text encoders, non-medical LDMs, or large data volumes. These strategies can limit performance and scientific accessibility. We propose a novel LDM conditioning approach to address these limitations. Methods: We propose Class-Conditioned Efficient Large Language model Adapter (CCELLA), a novel dual-head conditioning approach that simultaneously conditions the LDM U-Net with free-text clinical reports and radiology classification. We also propose a data-efficient LDM framework centered around CCELLA and a proposed joint loss function. We first evaluate our method on 3D prostate MRI against state-of-the-art. We then augment a downstream classifier model training dataset with synthetic images from our method. Results: Our method achieves a 3D FID score of 0.025 on a size-limited 3D prostate MRI dataset, significantly outperforming a recent foundation model with FID 0.071. When training a classifier for prostate cancer prediction, adding synthetic images generated by our method during training improves classifier accuracy from 69% to 74%. Training a classifier solely on our method’s synthetic images achieved comparable performance to training on real images alone. Conclusion: We show that our method improved both synthetic image quality and downstream classifier performance using limited data and minimal human annotation. Significance: The proposed CCELLA-centric framework enables radiology report and class-conditioned LDM training for high-quality medical image synthesis given limited data volume and human data annotation, improving LDM performance and scientific accessibility. Code from this study will be available at https://github.com/grabkeem/CCELLA
@@ 1 Introduction
Machine learning (ML) for medical image interpretation is a rapidly expanding field intent on improving imaging-based healthcare. However, recent ML advancements outside of medical imaging require training data in higher orders of magnitude than their predecessors. These advances cannot translate easily to the medical imaging space due to labeled data scarcity challenges including patient privacy and the cost of expert annotation. Recent medical imaging literature has seen an increase in latent diffusion model (LDM) studies [1]. Artificially increasing ML training volume with synthetic data produced by these generative methods provides a possible solution to labeled training data scarcity [2].
@@
Developing LDMs for medical imaging poses a few challenges. Training LDMs from scratch require large volumes of data. The Stable Diffusion LDM [3] for generating nonmedical, 2D color images was trained initially on 400 million image-text pairs [4]. A subsequent version of Stable Diffusion trained on 5.85 billion image-text pairs [5]. These data volumes are inaccessible to most healthcare institutions due to aforementioned data scarcity challenges. Non-medical 2D LDMs have been successfully fine-tuned on 2D medical images [6], however this strategy cannot easily extend to 3D. Adding labeled data conditioning has also shown to increase data efficiency during LDM training [7]. To this end, some medical LDMs have used the pre-trained CLIP text encoder to condition the LDM, similar to Stable Diffusion [3]. However, research has shown that the CLIP text encoder mostly emphasizes the first 20 tokens [8]. Thus, these medical LDM studies typically use only short, curated medical phrases or keywords with either the existing [9] or fine-tuned [6] CLIP text encoder rather than using the detail-rich clinical reports. These strategies require extra annotation and preprocessing or volumes of data that are challenging for a single institution to gather. This consequently limits the scientific accessibility of medical LDM development.
@@
In this study we propose Class-Conditioned Efficient Large Language Model Adapter (CCELLA), a text adapter extending ELLA [10], to improve LDM performance for synthetic medical image generation in a challenging medical imaging domain. CCELLA processes the text embedding generated by a large language model (LLM) from clinical reports prior to dual-head conditioning of the LDM U-Net. CCELLA performs two functions simultaneously: the first head aligns the text embedding with the image latent space prior to conditioning; the second head provides radiologist-derived class conditioning for the LDM during training. Each of these functions have independently demonstrated improvements in LDM performance outside of medical imaging [10, 11]. This joint function is hypothesized to provide two advantages: 1) to assert that the LDM receives conditioning about the presence or absence of suspicious lesions within the image, 2) more importantly, to inform CCELLA which text features should be emphasized when aligning the embedding to the image latent space for LDM conditioning. Notably, it is hypothesized that performing both functions in a single adapter will improve LDM performance beyond simply performing both functions independently. To achieve these functions, we also propose a modified loss function. This loss function simultaneously trains the LDM to produce high quality medical images and to appropriately extract radiology classification information from the clinical report text.
@@
Finally, we propose an LDM training framework around CCELLA that uses limited human data annotation and a training data volume reasonably accessible to a single healthcare institution. The intent of this framework is to improve the scientific accessibility of LDMs for medical imaging despite limited available data volumes. The framework functions by exploiting the clinical report and labeling already performed by radiologists during clinical routine and by reducing training requirements with publicly available, pre-trained components [2, 12].
@@
This study focused specifically on MRI for prostate cancer as the target domain to evaluate this framework. Prostate MRI is scored by radiologists using the Prostate Imaging – Reporting and Data System (PI-RADS) guidelines [13]. This pathology and imaging method was selected due to its difficulty in clinical interpretation, its moderate inter-reader variability between clinical experts, and the relatively low-availability of publicly available domain-specific data with the largest dataset consisting of 1500 samples [14, 15]. This study then performed a series of ablations to assess the mechanism by which CCELLA improves LDM performance and compared performance differences among variants of the proposed adapter. To validate utility of the proposed LDM training framework for dataset augmentation, this study demonstrated that the synthetic data generated with our proposed LDM framework can improve a classifier model at interpreting prostate MRI for cancer presence.
@@
  • A novel dual-head adapter for simultaneous alignment of medical text features extracted by a non-medical LLM to the medical imaging space and PI-RADS score extraction from medical text
  • A LDM training framework for improved radiology class-conditioned synthetic medical image generation that is trainable with a reasonable data volume.
@@ 2.1 Latent Diffusion Models
2.2 Conditioning and Medical Diffusion Models
LDMs have shown promising results in generating synthetic images both outside and within medical imaging [1]. Medical imaging LDM studies typically require either large data volumes to train from scratch, or strategic data-efficient methods to adapt pretrained LDMs. One critical, strategic approach is to condition the LDM with additional information (e.g. segmentation maps, text, or PI-RADS class) for improved image generation performance [18, 19, 2].
@@ 2.2 Conditioning and Medical Diffusion Models
Conditioning allows for control of the denoising U-Net behavior during LDM training and inference. LDM conditioning typically takes up to three forms. The first is timestep embedding, informing the LDM of its stage in the denoising process. Similar to positional encoding in transformers [20], the timestep is encoded with sinusoidal encoding and then conditions the LDM through addition with the latent image. The second is prompt embedding, informing the LDM of additional information during denoising. Prompt embedding can take different forms but typically are images or text. Prompting data are encoded with separate encoders and condition the LDM through concatenation with the latent image (usually for image prompting) or by cross-attention at multiple layers within the denoising U-Net (usually for text prompting). Class conditioning is a lesser-used third method that expands upon the timestep embedding by adding or concatenating a vector with class information to the timestep embedding [11]. This allows for class-based information to condition the U-Net during denoising. These conditioning methods can occur in parallel. For example, MAISI was trained to accept segmentation masks as prompt embedding, and included both image voxel spacing and body region as class conditioning that modified the timestep embedding [2].
@@
LDM text conditioning typically relies on the Contrastive Language-Image Pre-Training (CLIP) text encoder [21], trained to encode text to a shared image-text latent space. The CLIP text encoder is limited by its 77 text token input restriction, with literature documenting an effective length of under 20 tokens [22]. Consequently, text prompting for commercial LDMs are typically restricted to keywords or short phrases. This encoder was also not trained specifically for use with medical text. By contrast, clinical reports are typically longer and rich in free-text information (e.g. location, disease severity, or other defining characteristics) that could improve the realism of synthetic image generation [7]. Consequently, literature has developed multiple strategies for text prompting in medical imaging LDMs. These include crafting short phrases [23], training new embeddings for the frozen CLIP text encoder [18], fine-tuning the CLIP text encoder on medical data [6], or training new LDMs with a domain-relevant medical LLM [6, 19].
@@
PathLDM [7] is the most similar in motivation to our proposed work, as the authors of that study recognized the richness of clinical text, the value in using class-based information, the utility of automated ground-truth labeling, and the limitations to the CLIP text encoder for conditioning LDMs. PathLDM developed an automated text prompt generation workflow for synthetic 2D breast cancer histopathology slide generation. This workflow extracted two pathology-related probabilities (“Tumor” and “TIL”) with pretrained models that were each assigned “high” or “low” after probability thresholding, and summarized each slide’s clinical report to under 75 words using the GPT-3.5 LLM [25]. PathLDM also replaced the CLIP text encoder with a pretrained, histopathology-specific CLIP text encoder variant [26]. The final text prompt for LDM conditioning consisted of two class-based keywords and the summarized report in the form “[high/low] Tumor, [high/low] TIL, [report summary]”, allowing PathLDM to receive both classifier and key text report information as conditioning.
@@
  • Uses text from the original report instead of a shorter, LLM-generated summary;
  • Trains the class extractor simultaneously with the LDM, rather than relying on a non-medical, pretrained solution;
@@
  • Performs class conditioning using timestep embedding rather than through text encoding to enable class-specific training of the text adapter.
@@
  • Modifies each TSC block by adding PI-RADS score extractor layers;
@@ 3.1 Dataset
[…]
3.2 Latent Diffusion Model Implementation
The proposed LDM framework uses two pretrained components. The first is the MAISI LDM VQ-VAE, pretrained on 3D CT and MRI images [2]. The second is the FLAN-T5 XXL LLM [12]. The base LDM in this study used the frozen MAISI VQ-VAE with a learnable 4-channel, 4-layer denoising U-Net that concatenated image voxel spacings to the timestep embedding. The frozen FLAN-T5 XXL text encoder was used to encode the clinical report text.
CCELLA extends the ELLA adapter from [24] by adapting the encoded text for both deep cross-attention and timestep embedding conditioning. CCELLA begins with six class-conditioning TSC (CC-TSC) blocks, each consisting of a TSC block with a separate classifier head of two fully-connected layers of hidden size 256 and output size 2 (MLP block) at the end of the block. The outputs from all CC-TSC classifier heads were concatenated and fed into a final fully-connected layer to generate an output of size 2 to be concatenated with the timestep embedding. The image voxel spacing was additionally concatenated to the timestep embedding as per the foundation MAISI model [2]. The output of the final CC-TSC text embedding head conditioned the denoising U-Net through cross-attention. Figure 1 illustrates the CCELLA adapter design as well as the full dataflow for the LDM framework. Class conditioning was intentionally implemented as a separate vector for timestep embedding instead of through additional text prompt words as performed in PathLDM [7]. Our implementation enabled direct training of the CCELLA adapter using radiologist classification rather than relying on the adapter to train implicitly for improved image generation alone.
Figure 1: Overview of CCELLA-centric framework (left) and CCELLA adapter (right). Image encoder ($E_{Image}$) and decoder ($D_{Image}$) from [2]. Text encoder ($E_{Text}$) from [12]. CCELLA consists of six CC-TSC blocks with classifier ($h_{C}$) and aligned text embedding ($h_{T}$) heads. TSC block originally from [10]. Framework input includes the radiology report text for training and inference, with image and radiology classification additionally included for training alone. Example text redacted for privacy.
@@ 3.2 Latent Diffusion Model Implementation
The intent of this loss function is to emphasize both LDM image reconstruction as well as appropriate PI-RADS score extraction from the reference text. For this study, we used L1 loss for reconstruction and focal loss for class prediction, with class loss weighting $\lambda=10^{-4}$. L1 loss was selected due to its frequent use in LDM image reconstruction [3, 2]. Focal loss was selected for its success in the PI-CAI challenge [15]. Class loss weighting was determined empirically after training on a subsample of the full training set.
3.3 CCELLA Training and Evaluation
The MAISI and ELLA studies have already demonstrated success against state-of-the-art models [2, 24]. Specifically, the MAISI framework demonstrated improved Fréchet Inception Distance (FID) in all 2-dimensional planes compared to a pretrained, medical generative adversarial network [2]. ELLA demonstrated qualitative and quantitative improvement when adding their adapter to an existing commercial LDM [24]. Thus for this study, we assessed the CCELLA-centric framework against: 1) the MAISI LDM (MAISI) without text or class conditioning, 2) MAISI with radiology classification timestep conditioning (RADMAISI), 3) MAISI with text conditioning using ELLA to adapt FLAN-T5 XXL (ELLAMAISI).
Two ablation studies were then performed. The first assessed the components of CCELLA, namely the text and classifier conditioning heads as well as the joint loss function. This ablation study compared CCELLA to 1) an LDM with ELLA text conditioning and separate timestep conditioning using ground truth radiology classification (DECOUPLED), 2) CCELLA using only L1 loss instead of the proposed joint loss in (1) (L1ONLY). The second ablation study compared the CCELLA design to variants by: 1) placing only one CC-TSC block after a chain of five TSC blocks (LAST), 2) placing one CC-TSC block before five TSC blocks (FIRST), 3) chaining six CC-TSC blocks where class adapter weights were shared across all blocks (SHARED).
All models received image voxel spacing as timestep conditioning as per the foundation MAISI model [2].
Training details: All images were center-cropped to a $128mm\times 128mm$ field of view in the axial plane, normalized and then resampled to $256\times 256\times 128$ voxel dimensions using Lanczos interpolation [28] prior to LDM training and evaluation. All LDMs were trained for 400 epochs with initial learning rate of $10^{-4}$, the AdamW optimizer, a second-order polynomial learning rate scheduler, and batch size of 16 on four Tesla V100 GPUs. During the training of LDMs with text conditioning, images that did not have a text report used an empty string as a proxy text report. When training CCELLA, $\lambda$ in (1) was set to 0 for these cases so that the classifier portion of the loss function did not update the model for these images.
Evaluation: FID was used to evaluate LDM performances [29]. FID was calculated in two ways: 1) in 3-dimensions using the Med3D ResNet50 pretrained on 23 datasets [30], 2) in 2-dimensions for each of the sagittal, coronal and axial planes using a ResNet50 pretrained on RadImageNet [31].